# Few-shot LLM Synthetic Data with Distribution Matching

### v1

@@ A.2.2. Classifier Training Parameters
For the final classification task, we fine-tune a pre-trained language model on SST-2, AGNEWS, and Amazon datasets using the same set of hyperparameters. Specifically, we use a learning rate of 5e-5, a batch size of 32, and train for 6 epochs. A weight decay of 1e-4 is applied to prevent overfitting, and a warmup ratio of 6% is used to stabilize training by gradually increasing the learning rate during the initial steps.
@@ A.3. Prompt Design
To guide the synthetic data generation process, we design a two-stage prompting framework tailored to capture and generalize the linguistic attributes of each dataset. The full prompts for both stages are available in our code repository.
@@ A.3.1. Stage 1: Attribute Summarization
In the second stage, the LLM generates synthetic samples based on the summarized attributes from Stage 1. The prompt encourages diversity by instructing the model to create distinct examples that vary in subtopics and language styles. For instance, in the Amazon dataset, the LLM generates product reviews with varied focuses (e.g., different product features) while maintaining linguistic consistency. Similar instructions are adapted for the SST-2 and AGNEWS datasets to accommodate their specific attributes.

### v2

@@ A.2.2. Classifier Training Parameters
For the final classification task, we fine-tune a pre-trained language model using the parameters listed in Table 7. These include the learning rate (lr), batch size, training epochs, weight decay, and warmup ratio. The warmup ratio specifies the proportion of training steps used for learning rate warmup to stabilize training.
Table 7. Parameters for Classifier Training.
Parameter SST-2 AGNEWS Amazon
lr 5e-5 5e-5 5e-5
Batch size 32 32 32
Training epochs 6 6 6
Weight decay 1e-4 1e-4 1e-4
Warmup ratio 6% 6% 6%
A.2.3. Convex Hull Coverage
To calculate the coverage rate in 4.4.1, we first reduce the dimensionality of the sentence embeddings for both the original and generated datasets using t-SNE. A k-d tree is then constructed using the 2D t-SNE embeddings of the original dataset to enable efficient neighbor searching. The convex hull of the t-SNE embeddings of the original dataset is taken as the target distribution’s total coverage area. We initialize an empty buffer set $\mathcal{B}$ to store convex hulls formed during sampling. At each sampling iteration, a single example is selected, and its embedding is combined with the $k$-nearest neighbors (identified using the k-d tree) to create a new convex hull. If this convex hull overlaps with any existing convex hulls in $\mathcal{B}$, they are merged to form a larger convex hull. The total area of all convex hulls in $\mathcal{B}$ is then calculated and compared to the total coverage area of the original dataset to compute the coverage rate. This process is repeated iteratively for 200 sampling steps.
@@ A.3. Prompt Design
Figure 6. Prompt used in Stage 1 (Attribute Summarization). The LLM is instructed to analyze a dataset example and extract key linguistic and semantic attributes, such as topics, language habits, and writing styles, to guide the synthetic data generation process.
Figure 7. Prompt used in Stage 2 (Synthetic Data Generation). The LLM generates multiple unique samples based on the attributes summarized in Stage 1. For example, in the Amazon dataset, it generates reviews for similar products while ensuring diversity in subtopics and language styles.
To guide the synthetic data generation process, we design two stages of prompts tailored to capture and generalize the linguistic attributes of each dataset. The prompts are presented as examples in Figures 6 and 7.
@@ A.3.1. Stage 1: Attribute Summarization
In the first stage, the LLM analyzes a set of example data from the target dataset and generates a structured summary of its key linguistic and semantic attributes. The prompt instructs the model to extract attributes such as topics, language habits, and writing styles, depending on the dataset. For instance, for the SST-2 dataset, the attributes include movie genres, topics, language habits, and review length. This summarization step ensures that the generated synthetic data adheres to the original dataset’s characteristics.